find the exact value for the trigonometric expression?
tan(-pi/12)
HELP!
TY!

Answer by lwsshak3(11628)

find the exact value for the trigonometric expression?
tan(-pi/12)
**
Identity:tan(s-t)=(tan-tan t)/(1+tan s tan t)
tan(-pi/12)=tan(3π/12-4π/12)=tan(π/4-π/3)
tan(π/4-π/3)=(tan π/4-tan π/3)/(1+tan π/4*tan π/3)
=(1-√3)/(1+1*√3)=(1-√3)/(1+√3)
check:
using calculator
tan(-π/12)≈-0.2679..
(1-√3)/(1+√3)=-0.2679..